OpenStudy (anonymous):

Precalc question, please help? The point nearest to the origin on a line is at (4, -4). Find the standard form of the equation of the line.

OpenStudy (anonymous):

I know that the answer is x-y-8=0, I just don't know how to find that.

OpenStudy (anonymous):

The closest approach point (4, -4) lies on a line that is perpendicular to the line you have already. That means it has a slope which is -1/m compared with the slope m of the line you have. Get the new line's slope from change in y / change in x, from the (4,-4) point to the origin (0,0) (-4 - 0)/(4-0) = -1 The slope of your original line is thus -1/(-1) = m = 1 y = (1) x + b -4 = 4 + b so b = -8 y = x - 8, which is what you know the answer to be.

OpenStudy (anonymous):

Thank you!